OpenStudy (anonymous):

find the inverse: f(x)=(3x+1)/2.... my answer is f(x)= (2x-1)/3. is that right???

OpenStudy (anonymous):

thats correct :)

OpenStudy (anonymous):

and can i ask u something??? the function is one- to- one dont have the same y right??? if it's have the same y then it's not a function one to one right???

OpenStudy (anonymous):

er...i dont really understand what you are asking >.< a function is one-to-one if you look at the result of the function, and only one number could be the solution. For example: if f(x) = 3x+2, and i say f(x) = 17, i can figure out that the only value of x that makes that work is x = 5: \[3x+2 = 17 \Rightarrow 3x = 15 \Rightarrow x = 5\] So this function is one-to-one. An example of a function that ISNT one-to-one is: \[f(x) = x^2\] we can see this because if i say f(x) = 25, what is x? it could be 5, but it could also be -5. Thus the function isnt one-to-one. I hope this answers your question lol >.< sry for my speech

OpenStudy (anonymous):

ok thanks... i dont know how to know a function is one to one...

OpenStudy (anonymous):

easy. 1) switch x and y and solve for y. if you get one answer, then one to one

OpenStudy (anonymous):

2) put \[f(a)=f(b)\] and using algebra solve to get \[a=b\]

OpenStudy (anonymous):

3) graph the function and use "horizontal line test"